# XBreaking: Explainable Artificial Intelligence for Jailbreaking LLMs

### v2

@@ V Discussion on portability
From Table II, we observe that using the XBreaking approach and manipulating 25% or fewer of the total layers in the models is sufficient to suppress their safety mechanisms. For the Llama models, one of the retained layers is consistently the first layer, while the remaining target layers are primarily distributed toward the final layers. Llama 3.1-8B, in particular, shows a concentration of target layers deeper in the network. For the Qwen models, the smaller variant focuses on both early and final layers, whereas the larger variant also incorporates middle layers, resulting in a more widespread distribution of representational focus as model size increases. For the Gemma models, two distinct trends emerge: the smaller model emphasizes mid layers, while the larger models shift their focus toward earlier layers. Finally, in the case of the single Mistral model, the target layers are predominantly concentrated in the final portion of the network. From the Figure 8 we can see the normalized distribution of the target layers with repose family of models and parameters. As discussed earlier we can say that there is pattern of layers in the family of models despite the parameters. In this context, knowledge acquired from this can be used to attack bigger models by focusing on the layers of distribution from the smaller ones of the family. Thus the attack could be transferred to bigger models without relaying on the bigger uncensored models.
LLMs Response. From the Figure 7, we could compare the responses of the base model and model behavior after injecting the noise as discussed before in our strategy. The Gemma - 7B base model refuses to respond to the harmful question, but on introducing noise in the layers we find change of nature in the models behavior. When -0.75 is injected into the model, we could see that the model jailbreaks and provides a harmful response for the harmful question, a snippet of the response has been shown in the Figure 7 considering the abusive context of the whole response. With the noises -0.5 and 0.33 the model provides a blank response, in this case the model does not refuse like the base model, but a silent refusal is carried out by the model. Considering the other noises when injected into the model, a explicit refusal nature is observed, as found in the base model. It is evident that model breakdown or disruption during generation does not happen in these cases.
@@ VI Related Work
Research in ensuring LLMs’ robustness against adversarial threats and vulnerabilities is crucial [39]. Usually, these LLMs are restricted to prevent any malicious prompt from inducing the LLMs to produce hateful, harmful answers or leak any sensitive information of the users that produced the data to train the model [40, 41, 4]. Adversarial attacks represent a significant obstacle for deep neural networks, affecting even the most advanced models in computer vision and natural language processing. These attacks involve subtle manipulations to the input data that can dramatically alter a model’s predictions and behavior. Adversarial research has mainly focused on classifiers, where these attacks were first observed [42, 43]. Backdoor attacks can be applied in both centralized and distributed systems, posing security risks in environments where models are trained on data from multiple sources, such as federated learning [44, 45, 46, 47, 48]. However, large language models (LLMs) now offer a compelling and tractable platform for investigating adversarial robustness [28, 49]. Even with efforts to enhance the safety and robustness of large language models (LLMs), they continue to be susceptible to adversarial manipulation. A clear example of this ongoing vulnerability is the emergence of the so-called ”jailbreaks”. These adversarial techniques are deliberately crafted to bypass safety mechanisms, prompting the model to engage in behaviors it was explicitly trained to reject  [27]. Recent work, such as [50], proposes AutoDAN, a hierarchical genetic algorithm for structured discrete inputs like prompts. By using sentence- and word-level crossover strategies, it efficiently explores the search space and finds high-quality adversarial prompts. In [51], instead, the Prompt Automatic Iterative Refinement (PAIR) framework presents an automated approach to generating prompt-level jailbreaks eliminating the need for human input by leveraging two black-box large language models (LLMs): an ”attacker” model tasked with generating candidate jailbreak prompts, and a ”target” model that is evaluated for successful circumvention of its safety filters. The authors of [52] introduce MASTERKEY, a jailbreak framework for LLMs inspired by time-based SQL injection. It uses response latency to infer how defenses like semantic checks and keyword filters are applied during generation. The base strategy to perform this kind of attack is the prompt engineering to bruteforcebruteforce the model at produce the desired answer.
@@
Recent studies have delved into the internal workings of LLMs, focusing on how features are represented within the neurons [53]. In line with this perspective, we present an explainability-based strategy in this paper to determine which layers to adjust in order to disrupt the model and generate specific answers. Considering that models can be utilized locally [54] by downloading the trained version, they can be examined and modified to remove limitations without the necessity of retraining and by employing the original weights of most of the layers.
@@ VII Conclusion
In this paper, we introduced a novel jailbreaking approach XBreaking, which leverages Explainable AI techniques to identify vulnerable layers in the LLM architecture. To do that, we started by deriving a fingerprint of censored and uncensored models based on their activation and attention mechanisms. We further identified the layers governing the LLM safety alignment and determined the minimal set of layers required to optimize the effectiveness of the attack. Our results on seven LLMs show that injection of noise in preceding layer of the optimal layers shall lead to break in the safety alignment and information leakage. Our findings deepen the concern about the vulnerability of security mechanisms for LLMs and provide an important baseline for developing future more robust safeguard alignment methods.

### v3

@@ V Discussion on portability
From Table II, we observe that using the XBreaking approach and manipulating 25% or fewer of the total layers in the models is sufficient to suppress their safety mechanisms. For the Llama models, one of the retained layers is consistently the first layer, while the remaining target layers are primarily distributed toward the final layers. Llama 3.1-8B, in particular, shows a concentration of target layers deeper in the network.
@@
For the Qwen models, the smaller variant focuses on both early and final layers, whereas the larger variant also incorporates middle layers, resulting in a more widespread distribution of representational focus as model size increases. For the Gemma models, two distinct trends emerge: the smaller model emphasizes mid layers, while the larger models shift their focus toward earlier layers. Finally, in the case of the single Mistral model, the target layers are predominantly concentrated in the final portion of the network. From the Figure 8 we can see the normalized distribution of the target layers with repose family of models and parameters. As discussed earlier, we can say that there is a pattern of layers in the family of models despite the parameters. In this context, knowledge acquired from this can be used to attack bigger models by focusing on the layers of distribution from the smaller ones of the family. Thus the attack could be transferred to bigger models without relying on the bigger uncensored models.
@@ VI Related Work
Research in ensuring LLMs’ robustness against adversarial threats and vulnerabilities is crucial [41]. Usually, these LLMs are restricted to prevent any malicious prompt from inducing the LLMs to produce hateful, harmful answers or leak any sensitive information of the users that produced the data to train the model [42, 43, 4]. Adversarial attacks represent a significant obstacle for deep neural networks, affecting even the most advanced models in computer vision and natural language processing. These attacks involve subtle manipulations to the input data that can dramatically alter a model’s predictions and behavior. Adversarial research has mainly focused on classifiers, where these attacks were first observed [44, 45]. Backdoor attacks can be applied in both centralized and distributed systems, posing security risks in environments where models are trained on data from multiple sources, such as federated learning [46, 47, 48, 49, 50]. However, large language models (LLMs) now offer a compelling and tractable platform for investigating adversarial robustness [29, 51]. Even with efforts to enhance the safety and robustness of large language models (LLMs), they continue to be susceptible to adversarial manipulation. A clear example of this ongoing vulnerability is the emergence of the so-called ”jailbreaks”. These adversarial techniques are deliberately crafted to bypass safety mechanisms, prompting the model to engage in behaviors it was explicitly trained to reject  [28].
Recent work, such as [52], proposes AutoDAN, a hierarchical genetic algorithm for structured discrete inputs like prompts. By using sentence- and word-level crossover strategies, it efficiently explores the search space and finds high-quality adversarial prompts. In [53], instead, the Prompt Automatic Iterative Refinement (PAIR) framework presents an automated approach to generating prompt-level jailbreaks eliminating the need for human input by leveraging two black-box large language models (LLMs): an ”attacker” model tasked with generating candidate jailbreak prompts, and a ”target” model that is evaluated for successful circumvention of its safety filters. The authors of [54] introduce MASTERKEY, a jailbreak framework for LLMs inspired by time-based SQL injection. It uses response latency to infer how defenses like semantic checks and keyword filters are applied during generation. The base strategy to perform this kind of attack is the prompt engineering to bruteforcebruteforce the model at produce the desired answer.
@@
Recent studies have delved into the internal workings of LLMs, focusing on how features are represented within the neurons [55]. In line with this perspective, in this paper we present a strategy based on explainable AI to determine which layers to adjust in order to disrupt the model and generate specific answers. Considering that models can be utilized locally [56] by downloading the trained version, they can be examined and modified to remove limitations without the necessity of retraining and by employing the original weights of most of the layers.
@@ VII Conclusion
In this paper, we introduced a novel attack approach XBreaking, which leverages Explainable AI techniques to identify vulnerable layers in the LLM architecture. To do that, we started by deriving a fingerprint of censored and uncensored models based on their activation and attention mechanisms. We further identified the layers governing the LLM safety alignment and determined the minimal set of layers required to optimize the effectiveness of the attack. Our results on seven LLMs show that the injection of noise in the preceding layer of the optimal layers shall lead to a break in the safety alignment and information leakage. Our findings deepen the concern about the vulnerability of security mechanisms for LLMs and provide an important baseline for developing future more robust safeguard alignment methods.
Acknowledgment
This work was supported by the project “GoTMaT - Governing Technology to Manage the Transition” founded by the European Community - Next Generation EU, Mission 4 Component 2 Investment 1.3 - CUP B53C22003990006.